OpenStudy (anonymous):

If f(x) = x4 - 36, and g(x) = x2 + 6 then f(x) g(x) =

OpenStudy (anonymous):

f(x) = x4 - 36, and g(x) = x2 + 6 f(x)g(x)= (x^4 -36)( x^2 +6) = (x^4)(x^2)+6x^4 -36x^2 -36(6) = x^6 +6x^4 -36x^2-216

OpenStudy (anonymous):

(x4 - 36)(x2 + 6) x^6 - 6x^4 - 36x^2 - 216

OpenStudy (anonymous):

A) 2x3 - 3x2 - 50x - 75 B) 3x3 + x2 - 75x - 25 C) 6x2 - 7x - 3 D) 6x3 - 37x2 + 32x + 15 These are the answers.

OpenStudy (anonymous):

What?! wait what is the operation between the f(x) and g(x), multiplication right?

OpenStudy (anonymous):

Division

OpenStudy (anonymous):

If there's a division between f(x) and g(x) this should be the answer: .\[\LARGE {f(x)\over g(x)}={x^4-36 \over x^2+6}\] \[\LARGE ={x^4-6^2\over x^2+6}={(x^2-6)\cancel{(x^2+6)}\over \cancel{x^2+6} }=\] \[\LARGE =x^2-6 \]

OpenStudy (anonymous):

But that's not a answer on the question.

OpenStudy (anonymous):

then something's wrong, do you mind DRAWING your question ?

OpenStudy (anonymous):

Its sloppy.

OpenStudy (anonymous):

well, how can I help? give it a try...  this ? :O